# Hybrid Video Anomaly Detection for Anomalous Scenarios in Autonomous Driving

### v1

@@ 4 Evaluation
For the evaluation, we used the recently released AnoVox benchmark [17] with anomalous scenarios in the form of sudden braking maneuvers performed by lead vehicles, as visible in Figure 3. These are common and dangerous situations in road traffic [61] and are not included in our training data, as Figure 4 shows. In the 13 evaluation scenarios, the ego-vehicle follows a lead vehicle along a longer route under different weather and traffic conditions. At certain intervals, the lead vehicle performs a sudden braking maneuver. We extended the available pixel-wise ground truth, as shown in Figure 3 by frame-wise annotations. The annotation of the anomalies started with the unexpected sudden braking of the leading vehicle and ended with its complete stop.
Figure 4: Distribution of braking behavior in training (blue) and test (red) data. High-intensity sudden braking scenarios only occur in the evaluation dataset.
@@
In order to evaluate the performance of HF2-VADAD, we provide both quantitative and qualitative insights. Two metrics are used for evaluation: the already used Area Under the Receiver Operation Characteristic (AUROC) for frame-wise evaluations and the False Positive Rate at 95% True Positive Rate (FPR95) for pixel-wise evaluation. The AUROC metric, widely used in the VAD literature [62, 18, 63, 64]], serves as a measure of VAD accuracy by varying the threshold for anomaly assessment, with higher AUROC values indicating better detection performance. Adding FPR95 [65, 66] provides insights into the false-positive rates if nearly all anomalies are correctly detected, with lower values indicating better detection performance. Since our detected bounding boxes do not fully match the existing ground truth, we only evaluate the FPR95 based on the overlapping area and also provide the IoU to demonstrate the quality of the bounding boxes.
@@ 4.1 Quantitative Results
As shown in Table 2, we perform various experiments to better understand the performance of HF2-VADAD under varying conditions. Comparing city and highway scenarios, we observe a strong sensitivity of the model to these environments, but no clear trend emerges. For highway settings with bad weather, we can observe poor performance. As the pixel-wise anomaly scores are only calculated for bounding boxes and all other pixels are set to 0, there is generally a low false positive rate. However, if more than 5% of the ground truth anomalies are not detected, detecting all anomalous pixels requires a threshold close to zero. This results in the entire image being detected as anomalous, causing the FPR95 to rise sharply, as shown in Figure 5.
@@
To determine optimal values for the weights, as introduced in Section 3, we performed an empirical study, as shown in Table 3. While we observe the best results in the case when we completely neglect image predictions, we decided to set $w_{r}$, $w_{r_{p}}=10$ and $w_{p}$, $w_{p_{p}}=0.1$, as the performance metrics were too small to be significant, and we wanted to retain the advantages of both modules. The study showed that the model performs significantly worse when image prediction is weighted higher, which is in contrast to the original implementation of HF2-VAD, in which the frame prediction was weighted higher than the flow reconstruction, with the weight for frame prediction set to 1 and the weight for flow reconstruction set to 0.1.
Ablation Studies. To better understand the performance of our approach, we perform three types of ablation studies. First, we only utilize the intermediate outputs of the ML-ML-MemAE-SC to evaluate isolated flow reconstructions.
Second, we evaluate both bounding boxes generated from an off-the-shelf object detection model [60] and ground truth bounding boxes from the evaluation dataset to isolate the effect of model failures. We can observe that bounding boxes are best detected in simpler highway settings with good weather. The IoU performance suffers especially from poor detections of distant objects. Since mostly irrelevant vehicle are missing in the detections, we see improved FPR95 values for bounding box detections, but at the cost of reduced AUROC scores.
Lastly, we examine the influence of the fine-tuning stage during training. After fine-tuning, HF2-VADAD shows an improvement in precision. While the AUROC decreases slightly, the model achieves a better balance in the performance metrics with a significantly lower FPR95.
@@ 4.2 Qualitative Results
Figure 6: On the left, two normal scenarios are shown. Here, the model’s next-frame predictions (a) match the actual next frames (b) closely. On the right, we have shown sudden brake scenarios. Here, significant prediction differences arise, leading to high pixel-wise anomaly values (c).
In Figure 6, we show a series of comparative visualizations demonstrating the performance of HF2-VADAD in traffic scenarios. The model is able to predict regular behaviors very well, while atypical ones lead to poor predictions and, thus, high anomaly scores. The pixel-wise anomaly maps serve as a meaningful indicator, with more highlighted areas signaling larger prediction discrepancies. For normal driving patterns, these maps show only small deviations from ground truth, indicating that the model can predict the subsequent images well. For anomalies, the model often predicts, as it learned, that the vehicle maintains the distance. This leads to positive anomaly detections, as the vehicle is currently anomalous and getting closer due to the braking maneuver. This can be recognized by the increased highlighting in the difference maps.
Failure Cases. As shown in Figure 7, there are certain cases which are challenging. The first image shows how the model often incorrectly produces high anomaly scores when a lead vehicle is turning rather than going straight. The middle image shows a scene where the object detection model has problems with detecting objects in unfavorable weather conditions, such as heavy rain so anomaly detection is not possible.
Figure 7: Failure Cases: The model can produce high pixel-wise anomaly scores for turning vehicles, as shown on the left. Here, the encircled area in the frame-wise graph shows the duration of the curve. In the middle, a missing bounding box detection during bad weather situations leads to no anomaly detection at all. On the right, the bounding box only captures a part of the lead vehicle, and a truck in the back is completely missed.
@@ 5 Conclusion
In summary, our research has shown that HF2-VAD, a framework originally developed for detecting anomalies in surveillance systems can be effectively transferred to the field of autonomous driving. By extending it do dense pixel-wise anomaly maps and creating a dedicated training and test dataset, we have validated the effectiveness of HF2-VADAD in the vehicle context, bridging the gap between static monitoring and dynamic vehicle environments.
Our contributions further include the extension to dense anomaly detections, domain-specific weight adaptions, and dealing with multiple bounding boxes separately rather than combining them into one score, using robust scaling. In the future, we aim to further refine the framework to better deal with the known failure cases and evaluate more types of anomalies to examine its performance among a wider variety of scenarios, such as wrong-way drivers.

### v2

@@ 4 Evaluation
For the evaluation, we used the AnoVox benchmark [Bogdoll et al.(2024)Bogdoll, Hamdard, Rößler, Geisler, Bayram, Wang, Imhof, de Campos, Tabarov, Yang, Gottschalk, and Zöllner] with anomalous scenarios in the form of sudden braking maneuvers performed by lead vehicles, as visible in Figure 3. These are common and dangerous situations in road traffic [Kim et al.(2019b)Kim, Lee, Hwang, and Suh] and are not included in our training data, as Figure 4 shows. In the 13 evaluation scenarios, the ego-vehicle follows a lead vehicle along a longer route under different weather and traffic conditions. At certain intervals, the lead vehicle performs a sudden braking maneuver. Regularly behaving traffic participants are also present in the evaluation dataset. Based on the available pixel-wise ground truth annotations of sudden braking scenarios in AnoVox, as shown in Figure 3, we also generated bounding-box and frame-wise annotations for the evaluation. In AnoVox, the anomalies are annotated starting with the unexpected sudden braking of the leading vehicle and ending with its complete stop.
Figure 4: Distribution of braking behaviors of other traffic participants in training (blue) and test (red) data. High-intensity sudden braking scenarios, labeled as anomalous, only occur in the evaluation dataset.
@@
In order to evaluate the performance of HF2-VADAD, we provide both quantitative and qualitative insights. Two metrics are used for evaluation: the already used Area Under the Receiver Operation Characteristic (AUROC) for frame-wise evaluations and the False Positive Rate at 95% True Positive Rate (FPR95) for pixel-wise evaluation. The AUROC metric, widely used in the VAD literature [Gong et al.(2019)Gong, Liu, Le, Saha, Mansour, Venkatesh, and Hengel, Liu et al.(2021)Liu, Nie, Long, Zhang, and Li, Lu et al.(2013)Lu, Shi, and Jia, Mahadevan et al.(2010)Mahadevan, Li, Bhalodia, and Vasconcelos], serves as a measure of VAD accuracy by varying the threshold for anomaly assessment, with higher AUROC values indicating better detection performance. Adding the FPR95 [Henriksson et al.(2021)Henriksson, Berger, Borg, Tornberg, Sathyamoorthy, and Englund, Di Biase et al.(2021)Di Biase, Blum, Siegwart, and Cadena] metric provides insights into the false-positive rates if nearly all anomalies are correctly detected, with lower values indicating better detection performance. Since our detected bounding boxes do not fully match the existing ground truth, we only evaluate the FPR95 based on the overlapping area and also provide the IoU to demonstrate the quality of the bounding boxes.
@@ 4.1 Quantitative Results
As shown in Table 2, we perform various experiments with varying conditions. Comparing city and highway scenarios, we observe a strong sensitivity of the model to these environments, but no clear trend emerges. For highway settings with bad weather, we observe poor performance. As the pixel-wise anomaly scores are only calculated within bounding boxes and all other pixels are set to 0, there is generally a low false positive rate. However, if more than 5% of the ground truth anomalies are not detected, detecting all anomalous pixels requires a threshold close to zero. This results in the entire image being detected as anomalous, causing the FPR95 to rise sharply, as shown in Figure 5.
@@
To determine optimal values for the weights, as introduced in Section 3, we perform an empirical study, as shown in Table 3. While we observe best results in the case when we completely neglect image predictions, we decided to set $w_{r}$, $w_{r_{p}}=10$ and $w_{p}$, $w_{p_{p}}=0.1$, as the performance metrics were too small to be significant, and we retain the advantages of both modules. The study shows that the model performs significantly worse when image prediction is weighted higher, which is in contrast to the original implementation of HF2-VAD, in which the frame prediction was weighted higher than the flow reconstruction, with the weight for frame prediction set to $1$ and the weight for flow reconstruction set to $0.1$.
Ablation Studies. To better understand the performance of our approach, we perform three types of ablation studies, as shown in Table 2, with the focus on the first three entries in the lower part of the table. First, we only utilize the intermediate outputs of the ML-MemAE-SC to evaluate isolated flow reconstructions.
Second, we evaluate both bounding boxes generated from an off-the-shelf object detection model [MinyiLin(2022)] and ground truth bounding boxes from the evaluation dataset to isolate the effect of model failures. We observe that bounding boxes are best detected in highway settings with good weather. The IoU performance suffers especially from poor detections of distant objects. Since mostly irrelevant vehicles are missing in the detections, we see improved FPR95 values for bounding box detections, but at the cost of reduced AUROC scores.
Lastly, we examine the fine-tuning stage during training. After fine-tuning, HF2-VADAD shows an improvement in precision. While the AUROC decreases slightly, the model achieves a better balance in the performance metrics with a significantly lower FPR95.
@@ 4.2 Qualitative Results
Figure 6: On the left, two normal scenarios are shown. Here, the model’s next-frame predictions (a) match the actual next frames (b) closely. On the right, we have shown sudden brake scenarios. Here, significant prediction differences arise, leading to high anomaly values (c).
In Figure 6, we show a series of comparative visualizations demonstrating the performance of HF2-VADAD in traffic scenarios. The model is able to predict regular behaviors very well, while atypical ones lead to poor predictions and, thus, high anomaly scores. The pixel-wise anomaly maps serve as a meaningful indicator, with more highlighted areas signaling larger prediction discrepancies. For normal driving patterns, these maps show only small deviations from the ground truth, indicating that the model can predict the subsequent images well. For anomalies, the model often predicts, as it learned, that the vehicle maintains its distance. This leads to positive anomaly detections, as the vehicle is currently anomalous and getting closer due to the braking maneuver. This can be recognized by the increased highlighting in the difference maps.
Failure Cases. As shown in Figure 7, there are certain cases which are challenging. The first image shows how the model often incorrectly produces high anomaly scores when a lead vehicle is turning rather than going straight. This may be due to the model primarily learns straight-line driving patterns and the training data does not contain enough examples of lead vehicles turning directly in front of the car. The middle image shows a scene where the object detection model has problems with detecting objects in unfavorable weather conditions, such as heavy rain, so anomaly detection is not possible.
@@ 5 Conclusion
In summary, our research has shown that HF2-VAD, a framework originally developed for detecting anomalies in surveillance systems, can be effectively transferred to autonomous driving. By extending it to dense pixel-wise anomaly maps and creating a dedicated training and test dataset, we have validated the effectiveness of HF2-VADAD in the vehicle context, bridging the gap between static monitoring and dynamic vehicle environments.
Our contributions further include the extension to dense anomaly detections, domain-specific weight adaptions, and dealing with multiple bounding boxes separately rather than combining them into one score, using robust scaling. In the future, we aim to further refine the framework to better deal with current failure cases and evaluate more types of anomalies to examine its performance among a wider variety of scenarios, such as wrong-way drivers.
Figure 7: Failure Cases: Left: The model can produce high pixel-wise anomaly scores for turning vehicles. The encircled area in the frame-wise graph shows the duration of the curve. Middle: A missing bounding box detection during bad weather situations leads to no anomaly detection. Right: The bounding box only captures a part of the lead vehicle, and a truck in the back is completely missed.
Acknowledgment
This work results from the just better DATA project supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), grant number 19A23003H.